1. a. blue ray laser has a wavelength of 410 nm. a. what is the frequency...

1.

a.

blue ray laser has a wavelength of 410 nm. a. what is the frequency (HZ)

b. a 1-second flash of blue ray laser has 5.0E-3 J of energy. how many photons make up the flash ?

2. according to the Bohr model. if He+ ion emits a photon when an electron moves from n=4 to n=2 . will you be able to see this photon with your eyes? calculate the wavelength of the photon . recall visible light lies between 400 and 750 nm
Answer #1

1)

a)

Given:

lambda = 4.1*10^-7 m

use:

f = c/lambda

=(3.0*10^8 m/s)/(4.1*10^-7 m)

= 7.317*10^14 Hz

Answer: 7.32*10^14 Hz

b)

Given:

lambda = 4.1*10^-7 m

1st calculate energy of 1 photon

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(4.1*10^-7 m)

= 4.848*10^-19 J

number of photon = total energy/energy of 1 photon

n = 0.005/4.848*10^-19

= 1.031*10^16

Answer: 1.03*10^16

Only 1 question at a time please